# Kennedy-Tasaki transformation and non-invertible symmetry in lattice models beyond one dimension

### v1

@@ II.2 Kennedy-Tasaki transformation
Reference [25] showed that the KT transformation is equivalent to a sequence of operations: $TST$ or $STS$, where $S$ is gauging the global symmetry and $T$ is stacking a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT. Our definition here uses the picture of $S^{\dagger}TS$ that is similar to the picture of $STS$ up to a translation in its transformation. Our calculation verifies that $S^{\dagger}TS$ gives rise to the KT transformation. Explicitly, the action of our KT transformation is given by
@@
$\text{KT}^{\dagger}Z_{2i+1}=Z_{2i+1}\prod_{k=i+1}^{M}X_{2k}\,(\text{KT}^{\prime})^{\dagger},$ (22a)
$\text{KT}^{\dagger}Z_{2i}=\prod_{k=1}^{i-1}X_{2k+1}Z_{2i}\,(\text{KT}^{\prime\prime})^{\dagger}.$ (22b)
@@
which differs from Eq. (17) by the last part (without the Hermitian conjugation), and this will lead to additional translation. Furthermore, we have also verified that the alternative definition corresponding to $TST$ also give a valid KT transformation. We refer the readers to appendix B.
@@ II.3 Composition of operators
$\begin{split}\mathbf{D}_{SPT}&\equiv\text{KT}^{\dagger}\mathbf{D}\,\text{KT}\\ &=\mathbf{D}_{odd}^{\dagger}\mathbf{D}_{even}^{\dagger}\,T\,\mathbf{D}_{even}\mathbf{D}_{odd}\,T\,\mathbf{D}_{even}\mathbf{D}_{odd}\end{split}$ (25)
@@
On single $Z$ operators, $D_{SPT}$ act as
$\mathbf{D}_{SPT}Z_{2i+1}$ $=-Z_{1}Z_{2i+2}Y_{2M}\mathbf{D}_{SPT}^{\prime},$ (27a)
$\mathbf{D}_{SPT}Z_{2i}$ $=Y_{1}Z_{2}Z_{2i+1}\mathbf{D}_{SPT}^{\prime\prime},$ (27b)
@@
with $\mathbf{D}_{odd}^{\prime}$ and $\mathbf{D}_{even}^{\prime}$ given in Eq. (II.2) We remark that $\mathbf{D}_{SPT}$ squares to the projection and translation by two sites. Similar duality transformation can be achieved by $U_{CZ}\equiv\prod_{j}CZ_{j,j+1}$, although $(U_{CZ})^{2}=1$ and there is no translation.
@@ III.1 Two dimensions
As a first generalization, we will look at two dimensions. Recently, Ref. [36] provided a construction of a subsystem non-invertible symmetry operator that maps between self-dual models on different lattices. Our KW construction here builds upon the 1D result by Seiberg and Shao [16], which we have already reviewed in section. II.1.1; we map self-dual models on the same lattice.
@@
$\mathbf{P}$ $\equiv\prod_{j=1}^{L_{y}}\frac{(1+\eta^{x}_{j})}{2}\prod_{i=1}^{L_{x}}\frac{(1+\eta^{y}_{i})}{2}.$ (30c)
As it can be seen, $\tilde{\mathbf{D}}_{x}$ is composed of rows of the one-dimensional version of $\tilde{\mathbf{D}}$ in Eq. (2). Similarly, $\tilde{\mathbf{D}}_{y}$ is composed of columns of the one dimensional version of $\tilde{\mathbf{D}}$ in Eq. (2). $\mathbf{P}$ is a projector onto the subsystem symmetric subspace along both x and y directions. Now we define the following unitary action,
$\tilde{\mathbf{D}}^{(2)}\equiv\tilde{\mathbf{D}}_{x}\mathbf{H}\tilde{\mathbf{D}}_{y},$ (31)
where $\mathbf{H}$ represents the action of the Hadamard transformation on all sites (which transforms between $Z_{i,j}$ and $X_{i,j}$).
@@
$\tilde{\mathbf{D}}^{(2)}X_{L_{x},L_{y}}$ $=Z_{L_{x},L_{y}}Z_{1,L_{y}}Z_{L_{x},1}Z_{1,1}\eta^{x}_{L}\eta^{x}_{1}\tilde{\mathbf{D}}^{(2)}\eta^{y}_{L},$ (33c)
@@
$\tilde{\mathbf{D}}^{(2)}Z_{i,L_{y}}Z_{i,1}Z_{i+1,L_{y}}Z_{i+1,1}$ $=X_{i,1}\tilde{\mathbf{D}}^{(2)}\eta^{y}_{i}\eta^{y}_{i+1},$ (34b)
@@
$\mathbf{D}^{(2)}\equiv\mathbf{P}\tilde{\mathbf{D}}^{(2)}\mathbf{P}.$ (35)
@@
$\mathbf{D}^{(2)^{\prime}}\equiv\tilde{\mathbf{P}}^{\prime}\tilde{\mathbf{D}}^{(2)}\tilde{\mathbf{P}},$ (38)
@@
$\begin{split}\mathbf{\tilde{P}}^{\prime}&\equiv\prod_{l=1}^{j}\frac{\left(1-\eta^{x}_{l}\right)}{2}\prod_{l=j+1}^{L_{y}}\frac{\left(1+\eta^{x}_{l}\right)}{2}\prod_{k\neq 1,i,i+1}\frac{\left(1+\eta^{y}_{k}\right)}{2}\times\frac{(1-\eta^{y}_{i})}{2}\frac{(1-\eta^{y}_{i+1})}{2}\frac{(1+(-1)^{j}\eta^{y}_{1})}{2},\end{split}$ (39a)
$\mathbf{\tilde{P}}$ $\equiv\prod_{l\neq j}\frac{(1+\eta^{x}_{l})}{2}\prod_{k\neq i}\frac{(1+\eta^{y}_{k})}{2}\frac{(1-\eta^{x}_{j})}{2}\frac{(1-\eta^{y}_{i})}{2},$ (39b)
@@
$\begin{split}\mathbf{D}^{(2)\dagger}Z_{i,j}&=(-i)(-1)^{L_{x}-i+1}\prod_{k=i}^{L_{x}-1}\left(\left(\prod_{l=j}^{L_{y}-1}X_{k,l}\right)Y_{k,L_{y}}\right)\left(\prod_{l=j}^{L_{y}-1}X_{L_{x},l}\right)Y_{L_{x},L_{y}}Z_{L_{x},j}Z_{L_{x},j-1}\mathbf{\bar{P}}^{\prime}\mathbf{\tilde{\mathbf{D}}}^{(2)\dagger}\mathbf{\bar{P}},\end{split}$ (40)
@@
$\begin{split}\mathbf{\bar{P}}^{\prime}&\equiv\prod_{k=1}^{i-1}\frac{(1+\eta_{k}^{y})}{2}\prod_{k=i}^{L_{x}}\frac{(1-\eta_{k}^{y})}{2}\prod_{l\neq j,j-1,L_{y}}\frac{(1+\eta_{l}^{x})}{2}\frac{(1-\eta_{j}^{x})}{2}\frac{(1-\eta_{j-1}^{x})}{2}\frac{(1+(-1)^{L_{x}-i}\eta_{L_{y}}^{x})}{2},\end{split}$ (41a)
$\mathbf{\bar{P}}$ $\equiv\prod_{k\neq i}\frac{(1+\eta_{k}^{y})}{2}\prod_{l\neq j}\frac{(1+\eta_{l}^{x})}{2}\frac{(1-\eta_{k}^{y})}{2}\frac{(1-\eta_{l}^{x})}{2}.$ (41b)
Moreover, ater an appropriate phase choice for $\mathbf{D}^{(2)}$, $(\mathbf{D}^{(2)})^{2}$ is a translation $(i,j)\rightarrow(i+1,j+1)$ up to the projector:
$(\mathbf{D}^{(2)})^{2}=\mathbf{T}_{(1,1)}\mathbf{P},$ (42)
where $\mathbf{T}_{(1,1)}$ is a translation along the diagonal. The projection $\mathbf{P}$ is onto the subsystem symmetric Hilbert space compared to the projection in Eq. (8) that is onto 0-form symmetric subspace. Note that the translation in the diagonal direction leads to traversing of all sites if the lengths $L_{x}$ and $L_{y}$ have no common factor, i.e., $\text{gcd}(L_{x},L_{y})=1$.
@@ III.2 Higher dimensions
Let us consider a hypercubic lattice in $d$ spatial dimensions with coordinates $x_{1}$,…, $x_{d}$. Let us denote the set of all vertices of the hypercubic lattice by $\Delta_{v}$ and set of all hypercubes in the lattice by $\Delta_{c}$. We denote the boundary of a hypercubic cell $c$ by $\partial c$. The Hamiltonian for the TFHCI model is
@@
$\mathbf{D}^{(d)}\equiv\mathbf{P}\tilde{\mathbf{D}}_{d}\mathbf{H}\tilde{\mathbf{D}}_{d-1}\mathbf{H}\cdots\tilde{\mathbf{D}}_{1}\mathbf{P},$ (44)
where each of the $\tilde{\mathbf{D}}_{i}$ are string operators like Eq. (2) on straight line along $i$-th direction and $\mathbf{P}$ is a projection onto subsystem straight line-like symmetries. The operator $\mathbf{D}^{(d)}$ is a symmetry of the TFHCI model as it commutes with the Hamiltonian, i.e., $[\bm{H}_{\text{TFHCI}},\mathbf{D}^{(d)}]=0$. After an appropriate phase choice for $\mathbf{D}^{(d)}$, acting the operator twice will generate a diagonal unit shift in $d$ dimensions, i.e.,
$(\mathbf{D}^{(d)})^{2}=\mathbf{T}_{(1,...,1)}\mathbf{P}.$ (45)

### v2

@@ II.2 Kennedy-Tasaki transformation
Ref[27] showed that the KT transformation in the continuum is equivalent to a sequence of operations: $TST$ or $STS$, where $S$ is gauging the global symmetry and $T$ is stacking a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT. Our definition here uses the picture of $S^{\dagger}TS$ that is similar to the picture of $STS$ up to a translation in its transformation. Since we use a quantum circuit and projection to implement KW transformation, in our definition of $S$, we have a lattice translation by one unit. Therefore, $S^{\dagger}$ is not equal to $S$ on the lattice. Note that both $S^{\dagger}TS$ and $STS$ are the same in the continuum since in the continuum translation by one unit goes to identity operation and $S^{\dagger}$=$S$.
Our calculation verifies that $S^{\dagger}TS$ gives rise to the KT transformation. Explicitly, the action of our KT transformation is given by
@@
$\text{KT}^{\dagger}Z_{2i+1}=Z_{2i+1}\prod_{k=i+1}^{M{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-1}}X_{2k}\,(\text{KT}^{\prime})^{\dagger},$ (24a)
$\text{KT}^{\dagger}Z_{2i}=\prod_{k=1}^{i}X_{2k{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}1}Z_{2i}\,(\text{KT}^{\prime\prime})^{\dagger}.$ (24b)
@@
which differs from Eq. (19) by the last part (without the Hermitian conjugation), and this will lead to additional translation. Furthermore, we have also verified that the alternative definition corresponding to $TST$ also gives a valid KT transformation. We refer the readers to appendix B.
@@ II.3 Composition of operators
$\begin{split}\mathbf{D}_{SPT}&\equiv\text{KT}^{\dagger}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{D}_{even}\mathbf{D}_{odd}}\,\text{KT}\\ &=\mathbf{D}_{odd}^{\dagger}\mathbf{D}_{even}^{\dagger}\,T\,\mathbf{D}_{even}\mathbf{D}_{odd}\,T\,\mathbf{D}_{even}\mathbf{D}_{odd}\end{split}$ (27)
@@
On single $Z$ operators, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{D}}_{SPT}$ act as
$\mathbf{D}_{SPT}Z_{2i+1}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}Z_{1}Z_{2i+2}Y_{2M}\mathbf{D}_{SPT}^{\prime},$ (29a)
$\mathbf{D}_{SPT}Z_{2i}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}Y_{1}Z_{2}Z_{2i+1}\mathbf{D}_{SPT}^{\prime\prime},$ (29b)
@@
with $\mathbf{D}_{odd}^{\prime}$ and $\mathbf{D}_{even}^{\prime}$ given in Eq. (23). We remark that $\mathbf{D}_{SPT}$ squares to the projection and translation by two sites — or equivalently, by one site in each sublattice. Similar duality transformation as in Eq. (28) can be achieved by $U_{CZ}\equiv\prod_{j}CZ_{j,j+1}$ , which is the cluster entangler $T$ defined above, although $(U_{CZ})^{2}=1$ and there is no translation.
@@ III.1 Two dimensions
As a first generalization, we will look at two dimensions. Recently, Ref. [42] provided a construction of a subsystem non-invertible symmetry operator that maps between self-dual models on different lattices. Our KW construction here builds upon the 1D result by Seiberg and Shao [18], which we have already reviewed in the section. II.1.1; we map self-dual models on the same lattice.
@@
$\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}$ $\equiv\prod_{j=1}^{L_{y}}\frac{(1+\eta^{x}_{j})}{2}\prod_{i=1}^{L_{x}}\frac{(1+\eta^{y}_{i})}{2}.$ (32c)
As it can be seen, $\tilde{\mathbf{D}}_{x}$ is composed of rows of the one-dimensional version of $\tilde{\mathbf{D}}$ in Eq. (2). Similarly, $\tilde{\mathbf{D}}_{y}$ is composed of columns of the one-dimensional version of $\tilde{\mathbf{D}}$ in Eq. (2). $\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}$ is a projector onto the subsystem symmetric subspace along both x and y directions. Now we define the following unitary action,
$\tilde{\mathbf{D}}^{(2)}\equiv\tilde{\mathbf{D}}_{x}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}}\tilde{\mathbf{D}}_{y},$ (33)
where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}}$ represents the action of the Hadamard transformation on all sites (which transforms between $Z_{i,j}$ and $X_{i,j}$). We comment that the particular ordering of operators we set in defining Eq. (33) is our choice. Another possible definition is $\tilde{\mathbf{D}}^{(2)}\equiv\tilde{\mathbf{D}}_{y}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}}\tilde{\mathbf{D}}_{x}$. We note that the two definitions are related by swapping the $x$ and $y$ coordinates. Their action on Pauli operators $X_{i,j}$ and $Z_{i,j}$ operators are related by swapping $(i,j)\longleftrightarrow(j,i)$.
@@
$\tilde{\mathbf{D}}^{(2)}X_{L_{x},L_{y}}$ $=Z_{L_{x},L_{y}}Z_{1,L_{y}}Z_{L_{x},1}Z_{1,1}\eta^{x}_{L_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y}}}\eta^{x}_{1}\tilde{\mathbf{D}}^{(2)}\eta^{y}_{L_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x}}},$ (35c)
@@
$\tilde{\mathbf{D}}^{(2)}Z_{i,L_{y}}Z_{i,1}Z_{i+1,L_{y}}Z_{i+1,1}$ $=X_{i{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+1},1}\tilde{\mathbf{D}}^{(2)}\eta^{y}_{i}\eta^{y}_{i+1},$ (36b)
@@
$\mathbf{D}^{(2)}\equiv\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}\tilde{\mathbf{D}}^{(2)}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}.$ (37)
@@
$\mathbf{D}^{(2)^{\prime}}\equiv\tilde{\mathbf{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}^{\prime}}\tilde{\mathbf{D}}^{(2)}\tilde{\mathbf{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}},$ (40)
@@
$\begin{split}\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}^{\prime}}&\equiv\prod_{l=1}^{j}\frac{\left(1-\eta^{x}_{l}\right)}{2}\prod_{l=j+1}^{L_{y}}\frac{\left(1+\eta^{x}_{l}\right)}{2}\prod_{k\neq 1,i,i+1}\frac{\left(1+\eta^{y}_{k}\right)}{2}\times\frac{(1-\eta^{y}_{i})}{2}\frac{(1-\eta^{y}_{i+1})}{2}\frac{(1+(-1)^{j}\eta^{y}_{1})}{2},\end{split}$ (41a)
$\mathbf{\tilde{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}$ $\equiv\prod_{l\neq j}\frac{(1+\eta^{x}_{l})}{2}\prod_{k\neq i}\frac{(1+\eta^{y}_{k})}{2}\frac{(1-\eta^{x}_{j})}{2}\frac{(1-\eta^{y}_{i})}{2},$ (41b)
@@
$\begin{split}\mathbf{D}^{(2)\dagger}Z_{i,j}&=(-i)(-1)^{L_{x}-i+1}\prod_{k=i}^{L_{x}-1}\left(\left(\prod_{l=j}^{L_{y}-1}X_{k,l}\right)Y_{k,L_{y}}\right)\left(\prod_{l=j}^{L_{y}-1}X_{L_{x},l}\right)Y_{L_{x},L_{y}}Z_{L_{x},j}Z_{L_{x},j-1}\mathbf{\bar{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}^{\prime}}\mathbf{\tilde{\mathbf{D}}}^{(2)\dagger}\mathbf{\bar{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}},\end{split}$ (42)
@@
$\begin{split}\mathbf{\bar{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}^{\prime}}&\equiv\prod_{k=1}^{i-1}\frac{(1+\eta_{k}^{y})}{2}\prod_{k=i}^{L_{x}}\frac{(1-\eta_{k}^{y})}{2}\prod_{l\neq j,j-1,L_{y}}\frac{(1+\eta_{l}^{x})}{2}\frac{(1-\eta_{j}^{x})}{2}\frac{(1-\eta_{j-1}^{x})}{2}\frac{(1+(-1)^{L_{x}-i{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+1}}\eta_{L_{y}}^{x})}{2},\end{split}$ (43a)
$\mathbf{\bar{P}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}$ $\equiv\prod_{k\neq i}\frac{(1+\eta_{k}^{y})}{2}\prod_{l\neq j}\frac{(1+\eta_{l}^{x})}{2}\frac{(1-\eta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}}^{y})}{2}\frac{(1-\eta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}j}}^{x})}{2}.$ (43b)
Moreover, after an appropriate choice of normalization for $\mathbf{D}^{(2)}$, $(\mathbf{D}^{(2)})^{2}$ is a translation $(i,j)\rightarrow(i+1,j+1)$ up to the projector:
$(\mathbf{D}^{(2)})^{2}\propto\,\mathbf{T}_{(1,1)}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}},$ (44)
where $\mathbf{T}_{(1,1)}$ is a translation along the diagonal. The projection $\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2)}}$ is onto the subsystem symmetric Hilbert space compared to the projection in Eq. (8) that is onto a 0-form symmetric subspace. Note that the translation in the diagonal direction leads to traversing of all sites if the lengths $L_{x}$ and $L_{y}$ have no common factor, i.e., $\text{gcd}(L_{x},L_{y})=1$.
We note that the KW transformation can also be implemented by unitary gates plus measurements in a short-depth operation, which maps the degrees of freedom to those on the dual lattice; see Appendix F.
@@ III.2 Higher dimensions
Let us consider a hypercubic lattice in $d$ spatial dimensions with coordinates $x_{1}$,…, $x_{d}$. Let us denote the set of all vertices of the hypercubic lattice by $\Delta_{v}$ and the set of all hypercubes in the lattice by $\Delta_{c}$. We denote the boundary of a hypercubic cell $c$ by $\partial c$. The Hamiltonian for the TFHCI model is
@@
$\mathbf{D}^{(d)}\equiv\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(d)}}\tilde{\mathbf{D}}_{d}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(d)}}\tilde{\mathbf{D}}_{d-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(d)}}\cdots\tilde{\mathbf{D}}_{1}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(d)}},$ (46)
where each of the $\tilde{\mathbf{D}}_{i}$ are string operators like Eq. (2) on a straight line along $i$-th direction , ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(d)}}$ represents the tensor product of Hadamard operators on all the sites of $d$ dimensional lattice and $\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(d)}}$ is a projection onto subsystem straight line-like symmetries. The operator $\mathbf{D}^{(d)}$ is a symmetry of the TFHCI model as it commutes with the Hamiltonian, i.e., $[\bm{H}_{\text{TFHCI}},\mathbf{D}^{(d)}]=0$. After an appropriate choice of normalization for $\mathbf{D}^{(d)}$, acting the operator twice will generate a diagonal unit shift in $d$ dimensions, i.e.,
$(\mathbf{D}^{(d)})^{2}\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\propto}\,\mathbf{T}_{(1,...,1)}\mathbf{P}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(d)}}.$ (47)